OpenStudy (ttp):

Which number is not in scientific notation? one point one times ten to the power of two hundred thirteenth seven point eight times ten to the eighth power three point zero two times ten to the negative tenth power thirty-five times ten to the fourth power

OpenStudy (anonymous):

"one point one times ten to the two hundred thirteenth" \[1.1 \times 10^{200/13}\] "seven point eight times ten to the eighth" \[7.8 \times 10^8\] three point zero two times ten to the negative tenth \[3.02 \times 10^{-8}\] thirty-five times ten to the fourth \[35 \times 10^4\]

OpenStudy (anonymous):

D) is not in scientific notation because it has 2 in front of decimal point I'm not sure about A)

OpenStudy (ttp):

I still don't know which one the answer is?

OpenStudy (anonymous):

200/13 = 15.384 10^15.384 ? I don't know if thats allowed in scientific notation....

OpenStudy (ttp):

Mhm so D or A?

OpenStudy (anonymous):

only one choice allowed right ?

OpenStudy (ttp):

Yesum

OpenStudy (ttp):

It's only one answer ...

OpenStudy (anonymous):

in that case A) is violating rules

OpenStudy (anonymous):

D) is not normalized but at least it's not violating the rules

OpenStudy (anonymous):

(a times ten raised to the power of b), where the exponent b is an integer, and the coefficient a is any real number (however, see normalized notation

OpenStudy (ttp):

Ok thanks. !